OpenStudy (anonymous):

simplify this radical expression please? sqrt 3+ 4 sqrt 3. Possible answers: 5sqrt 6, 5sqrt 3, 3 sqrt 6, or 3 sqrt 3.

OpenStudy (sburchette):

You can think of sqrt 3 and 4sqrt 3 as being like terms. All you have to do is add the coefficients of the radicals.\[\sqrt{3}+4\sqrt{3}=(1+4)\sqrt{3}=5\sqrt{3}\]

OpenStudy (anonymous):

Where did that 1 come from?

OpenStudy (sburchette):

The one is an implied coefficient of sqrt3. Similar to how x is also 1x.